b3lla2025:

a certain kind of lumber costs $3.00 for 8 ft. at this rate how much does a piece that is 14 ft long cost

b3lla2025:

@supie

b3lla2025:

help

XxThomasxX:

3/8=x/14 8x=42 x=42/8 or 5.25 (5 1/4) (21/4)

b3lla2025:

I got 24

XxThomasxX:

If you buy $3 for $8, the money is less than the wood. If you buy 24 for 14, that is more. just try 5.25 or however u do it.

b3lla2025:

what

b3lla2025:

@supie help please

XxThomasxX:

just try it

XxThomasxX:

its correct

XxThomasxX:

3/8=x/14. Cross multiply. 8*x=3*14 8x=42. x=42/8 or 5.25

b3lla2025:

so 5 dollars and 25 cents

XxThomasxX:

yup

b3lla2025:

thxs

XxThomasxX:

yw always

b3lla2025:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XxThomasxX yw always \(\color{#0cbb34}{\text{End of Quote}}\) can you help me with one more question

XxThomasxX:

sure

b3lla2025:

a package of 3 lb of ground beef costs $5.10. how much ground beef could you buy from $3.40 at this rate

XxThomasxX:

3/5.10=x/3.40 5.10x=10.2 x=1.02 so you could buy 1.02 pounds of ground beef

b3lla2025:

thxs

XxThomasxX:

yw anytime